-> Jun 12, 2020 DOMC experienced a problem of interference with another experiment at DomeC, starting 08-Jun-2020 ca. 14 UT. Repair work is in process.
-> Jul 26, 2019 Lead-free neutron monitor DOMB is out of service (upgrade of electronics).
-> Jun 11, 2014 Section A1 was replaced with a spare one. A small change of the count rate is possible, will be corrected later.
-> Mar 14, 2014 The data server computer has been replaced. No change to the measuring computers => no data lost.
-> Feb 10, 2012 A problem in the main registration channel 10:51-12:48 UT 09-Feb-2012. Data from the redundant channel were used.
-> Apr 14, 2011 Power break occurred on 13-Apr-2011. Data lost for 02:27-02:42 UT.
-> Oct 29, 2010 Problems with Internet connections occurred on 28-29-Oct-2010. No data lost.
-> July 26, 2010 Prophylaxis and maintenance work is in progress. Some sections can be OFF. Data statistics can be worsen.
-> December 17, 2009 Prophylaxis work took place on December 16, 2009 between 07:48 and 13:00 UT. One section was off consequently (A,B, and then C), data used from the two remaining sections.
-> November 11, 2009 The new registration system was off from 16:53 UT 2009.11.07 till 12:50 UT 2009.11.09 because of a software problem. Data from the redundant old registration system were used during this period.
-> February 05, 2009 New Linux-based data server has been launched. The old server is available as http://oldcr.oulu.fi. The old server is not maintained any longer. The data there are "as is", i.e. not verified.
-> January 07, 2009 Due to a software problem, the new registration system was off during 01-03 January 2009. Data have been recorded by the old system. No data have been lost. The new system is fully operational again since 09:19 UT 03-Jan-2009.
-> November 10, 2008 Power break resulted in a data gap 10:55-11:25 UT 10-Oct-2008.
-> August 22, 2008
New registration system is writing data to the data-base since August 1st, 2008. Data are updated on the server every minute. Because of slightly higher efficiency of the new system, the new normalization factor needs to be applied for long-term studies.
-> November 19, 2007
A new data recording system is under test. Some data for the last few days might be affected. All data are corrected and reduced to the standard counts. 10-sec channel is switched off.
-> August 13, 2007
Collecting computer crashed at 07:00 05-Aug-2007 and restarted at 11:05 08-Jul-2007. Data restored using 10-sec channel, pressure obtained from VTT Oulu.
-> July 04, 2007
Collecting computer crashed at 21:00 03-Jul-2007 and restarted at 06:00 04-Jul-2007. Data restored using 10-sec channel, pressure obtained by interpolation.
-> Feb 06, 2007
The primary detecting channel was off 00:00 UT 05/02/2007 - 06:52 UT 05/02/2007 because of the GPS error. Data have been recorded by the secondary channel.
-> Jan 22, 2007
The primary detecting channel was off 15:00 UT 21/01/2007 - 07:00 UT 22/01/2007 because of a GPS error. Data have been recorded by the secondary channel.
-> Dec 13, 2006
A strong GLE has occurred on 13 Dec 2006, >90% increase in 5-min data.
-> Sept 04, 2006
02-04 Sept. 2006 : Intermittent software problem in the collecting computer. Data restored using 10-sec channel, pressure obtained from FMI and interpolated.
-> Aug 30, 2006
29-30 Aug. 2006 : Collecting computer software failure. Data restored using 10-sec channel, pressure obtained from FMI.
-> Aug 22, 2006
20-22 Aug. 2006 : Problems with electric power for the server. Data collecting front end computer uses UPS => No data have been lost.
-> Jan 02, 2005
The list of publications based on the use of Oulu NM data has been updated and counts 94 items now.
-> Jul 06, 2005
The station server crashed on 02 July 2005. It was fully recovered only on 05 July 2005. Due to software installation work, pressure data were lost for some periods during 10-14 UT 05 Jul. 2005. Interpolations were used to reconstruct the pressure data. No cosmic ray data have been lost.
-> Dec 24, 2003
5-minute resolution data for Jun - Oct 1984 have been added to the database. The 5-minute database is completed!
-> Dec 17, 2003
All counters are working.
-> Nov 02, 2003
Section A is off since 07:30 UT 08 Dec 2003 for the purpose of hardware reconstruction.
-> Nov 02, 2003
Two GLE's (Nos.65 and 66) have been added to the GLE list
-> Oct 15, 2003
5-minute resolution data for Jan - Jun 1977 have been added to the database
-> Oct 10, 2003
A problem with Section A has been fixed for two counter tubes. Section A is on since 10 Oct. 2003, 14 UT. The third counter tube of Section A is not repaired yet.
-> Oct 08, 2003
The problem with Section B has been fixed, two sections are in operation since 09 UT 08 Oct. 2003.
-> Oct 06, 2003
Beacuse of hardware problems, only one section (C) is in operation.
-> Oct 01, 2003
5-minute resolution data for Jul - Dec 1976 have been added to the database
-> Sep 25, 2003
5-minute resolution data for Jan - May 1984 have been added to the database
-> Sep 24, 2003
5-minute resolution data for Jan - Jun 1976 have been added to the database
-> Sep 15, 2003
5-minute resolution data for Jul - Dec 1975 have been added to the database
-> Sep 02, 2003
5-minute resolution data for Jul - Dec 1983 have been added to the database
-> Sep 02, 2003
Correction coefficient has been changed since June 2003 because of the Section A high voltage adjustment - see here
-> Sep 01, 2003
5-minute resolution data for Jan - Jun 1983 have been added to the database
-> Aug 29, 2003
5-minute resolution data for Jan - Dec 1982 have been added to the database
-> Aug 20, 2003
5-minute resolution data for Jan - Dec 1981 have been added to the database
-> Aug 19, 2003
5-minute resolution data for Jan - Dec 1980 have been added to the database
-> Aug 15, 2003
5-minute resolution data for Jul - Dec 1979 have been added to the database
-> Jul 27, 2003
5-minute resolution data for Jan - Jun 1979 have been added to the database
-> Jul 04, 2003
5-minute resolution data for July-December 1977 have been added to the database
-> Jun 02, 2003
Power break on 17 June 2003 - data are lost between 17:00 and 19:15 UT
-> Jun 02, 2003
The IP adress has been changed 184.108.40.206 -> 220.127.116.11.
-> February 28, 2003
The web server has been recovered. No data have been lost.
-> February 26, 2003
The web server has crashed. The system was re-installed.
-> February 20, 2003
Section B was off for profilactics on 20 February 2003 between 13:47 and 15:00 UT. Data from Sections A and C were used during that period.
-> February 10, 2003
Pressure values have been corrected for 24 Nov 2001.
-> December 20, 2002
The high voltage supplier of Section C (3 NM64) is fixed. The section is on since about 13:00 UT of December 20, 2002.
-> December 19, 2002
Section C (3 NM64) has been switched off at 13:30 UT on Dec. 19, 2002 because of a high voltage supplier failure. We hope to fix the problem soon.
-> October 29, 2002
5-min data for Jan-Jun 1975 have been digitized and added to the database.
-> September 23, 2002
5-min data for Jun-Dec 1974 have been digitized and added to the database.
-> Sept. 02, 2002
Data of GLE # 64 on 24 August 2002 are added.
-> July 22, 2002
5-min data for Jan-May 1974 have been digitized and added to the database.
-> July 01, 2002
5-min data for Jul-Dec 1973 have been digitized and added to the database.
-> May 21, 2002
5-min data for Jan-Jun 1973 have been digitized and added to the database.
-> April 22, 2002
Section C was off (for profilactics reason) during 08:30 - 12:15 UT on 22 Apr. 2002.
-> February 04, 2002
5-min data for 01-06 / 1972 are digitized and added to the database.
-> January 15, 2002
5-min data for 1971 and 07-11/1972 are digitized and added to the database.
-> January 09, 2002
Data of GLE # 63 on 26 December 2001 are added.
-> January 09, 2002
The GPS system was not working during 23.Dec.2001-09.Jan.2002. The timing during that period was within few minutes accuracy. On 09.Jan.2002 GPS is fixed - timing is within 1 sec now.
-> November 12, 2001
Data of GLE # 62 on 04 November 2001 are added.
-> October 31, 2001
We started to digitize earlier 5-min data which were only available in the printed paper form. Now 5-min data are available for July 1968 - December 1970. We keep working digitizing the data and expanding the interval of restored 5-min early data.
-> October 10, 2001
All data gaps are filled for 1990-1991 using the restored archive of 10-sec data.
-> August 16, 2001
The list of publications using data of Oulu NM is updated.
-> July 24, 2001
The server was down on 23/07/2001 because of an Internet-worm attack. Data collection was OK. The problem is fixed, data are NOT infected.
-> July 20, 2001
Due to reconfiguration of GPS, data for 3 minutes, 09:27 - 09:29 UT 20/07/2001, were lost.
-> May 14, 2001
The GPS is fixed. The timing is precise again.
-> May 09, 2001
GLE-formatted data files are added for the two April 2001 GLE's
-> April 19, 2001
The GPS was broken during the magnetic storm. Until it is fixed, the timing is only within 15-sec accuracy.
-> March 26, 2001
The above mentioned bug is fixed. If You downloaded digital data before 26.03.2001, it is recommended that You download them again.
-> March 24, 2001
A bug in the program has been found (thanks to Roger Pyle). Digital data (in ASCII format) were incorrectly scaled for some periods. Graphical output is correct.
-> March 20, 2001
Links to the data in GLE format are added.
-> February 2001
The data base is completed. Data since 1964 are included with the best possible resolution: hourly data since 1964, 5-min since 1985, 1-min since 1996.
-> November 2000
The new server has been launched with data since 1996. Work of Boris Gvozdevsky (PGI, Apaptity) is greatly acknowledged.
Back to the main page
Last modified 26.03.2001 by Ilya